Helped by a strong jet stream, a Los Angeles-to-Boston plane flew 10% faster than usual and made the 4400 km trip in 30 min less time than usual. At what speed does the plane usually fly?
I tried a chart but I don't think I set it up right.

Answer by ewatrrr(24785)


Hi,
Let x represent the speed the plane usually flies
d = r*t
4400km = x*t
t = 4400/x
4400km = (x + .10x)(t -.5hr)
4400km = 1.10x(4400/x -.5hr)
4400km =4840-.55hr*x
.55x = 440
x = 800km/hr